Shift Summary for Day Shift, Jan. 17, 00


Subject: Shift Summary for Day Shift, Jan. 17, 00
From: Vivian O'Dell (odell@fnal.gov)
Date: Fri Feb 04 2000 - 15:28:54 CST


         Shift Summary for Day Shift, Jan. 17, 00

On Shift: Brad C., M. Arenton

          plus cast of thousands come to sample the
          champagne

Hours of beam available at all intensities: 0.0 hours
Hours of beam available at nominal intensity: 0.0 hours
Hours of E799 data (special runs not included): 0.0 hours
Hours of special runs (laser scans, muon runs...): 1.0 hours

Accounting of lost time:

End of E799 run. Last beam at ~8:01 AM
Change over to KAMI test

========================================================================
Runs ended this shift:
  
none

=======================================================================
Tapes taken to FCC this shift

UPA331-333 UPS111
UPB331-333
UPC331-333

UPH059, UPK042, UPF079

========================================================================
Accesses during shift:

Area open

========================================================================
Overview of shift:

8:00 Analysis magnet off; Toasts to the end of E799 run
8:05 Turn off BA, HG high voltages, request turn off of NM3S
      Interlock dropped
8:30 long laser run started
8:30 BA uncabling underway
8:40 Survey crews arrive, lock out DC HV and prepare to survey
      Drift chambers.
8:45 NM3 is opened and gate valve closed and locked. Crew
      arrives to survey and swap defining collimator insert.
9:20 BA uncabeling is finished.
9:30 Troubles with laser scan traced to a clear that had not
      been established when using a fake spill. Now running
      laser scan.
10:00 Concrete block removal begins.
10:30 Good laser scan run 15553 finished, data to disk only
10:50 Tapes taken to Feynman.
11:40 Concrete blocks are all out. Instead of the 5 large blocks
      expected, there were actually 15 smaller ones.
13:25 Old BA and stand removed.
14:00 New BA stand dropped in
14:15 BAH dropped into place
14:35 BAG dropped into place
15:00 Replacing shielding blocks
15:30 shift ends

========================================================================
Description of Shift:
 
"controlled chaos", lots of work

========================================================================
Follow up Notes on Previous Shift

Continue with KAMI setup



This archive was generated by hypermail 2b27 : Fri Feb 04 2000 - 15:29:01 CST